OpenStudy (anonymous):

which is an equation in point-slope form for the given point and slope point (-3 7) slope 4 A y-7=4(x-3) B y-7=4(x+3) C 7+7=4(x+3) Dy+7=4x-12

OpenStudy (welshfella):

this form is y - y1 = m(x - x1) where m = slope and (x1,y1) is the given point

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok so what is the y-y1

OpenStudy (welshfella):

well (x1,y1) = (-3,7) so x1 = -3 and y1 = 7

OpenStudy (welshfella):

you just replace m , x1 and y1 in the equation by the given values

OpenStudy (welshfella):

ill start you off y - y1 = m(x - x1) y - 7 = can you do the right side?

OpenStudy (welshfella):

m = 4 and x1 = -3

OpenStudy (anonymous):

oh ok i think i get it , is the answer y-7=4(x-3)

OpenStudy (welshfella):

nearly right its x + 3 in the panertheses beacuse x1 = -3 y - 7 = 4(x - (-3)) - note the double negative - this simplifies to a '+' so its y - 7 = 4(x + 3)

OpenStudy (welshfella):

-(-3) = + 3 ok?

OpenStudy (anonymous):

ok thank you for your help